# GigaBrain-0: A World Model-Powered Vision-Language- Action Model

### v1

@@ 4.2 World-Model-Generated Data
Generated Data Quality Inspection. Generated videos inevitably contain hallucinations or artifacts that may degrade downstream training performance. To mitigate this, GigaWorld introduces a comprehensive quality assessment pipeline that evaluates generated videos across multiple dimensions: geometric consistency (Liu et al., 2025a), multiview consistency (Liu et al., 2025a), text-description alignment (Azzolini et al., 2025), and physical plausibility (Azzolini et al., 2025). Each video is assigned a composite quality score, which determines whether it is suitable for pre-training, fine-tuning, or should be discarded. All model and training details mentioned above will be fully elaborated in the upcoming GigaWorld technical report.
@@ 5.1 Dexterous Manipulation Experiment
Experiment Setup. We evaluate GigaBrain-0 on two dexterous manipulation tasks, laundry folding and paper towel preparation. For laundry folding, we deploy the model on the G1 robot platform and fine-tune it using 300 human-collected demonstration trajectories, the training is conducted with a batch size of 128 for 40K steps. For paper towel preparation, we use 100 demonstrations from PiPER arms, and the training is conducted with a batch size of 128 for 20K steps. We compare GigaBrain-0 against $\pi_{0}$ (Black et al., 2024), which is implemented using the official open-source code, and fine-tuned with the same training config for fair comparison.
@@ 5.2 Long-horizon Experiment
Experiment Setup. We evaluate GigaBrain-0 on two long-horizon manipulation tasks: table bussing and juice preparation. For table bussing, we deploy the model on the dual-arm PiPER robot platform and fine-tune it using 100 human-collected demonstration trajectories, and the training is conducted with a batch size of 128 for 20K steps. For juice preparation, we utilize the AgiBot G1 robot with 489 demonstrations, and the training is conducted with a batch size of 128 for 35K steps. We compare GigaBrain-0 against $\pi_{0}$ (Black et al., 2024), which is implemented using the official open-source code, and fine-tuned with the same training config to ensure a fair comparison.
@@ 5.3 Mobile Manipulation Experiment
Experiment Setup. We evaluate GigaBrain-0 on two mobile manipulation tasks: boxes moving and laundry baskets moving. For boxes moving, we deploy the model on the AgiBot G1 robot platform and fine-tune it using 300 human-collected demonstration trajectories, and the training is conducted with a batch size of 128 for 30K steps. For laundry baskets moving, we utilize the dual-arm PiPER robot platform with 378 demonstrations, and the training is conducted with a batch size of 192 for 30K steps. We compare GigaBrain-0 against $\pi_{0}$ (Black et al., 2024), which is implemented using the official open-source code, and fine-tuned with the same training config to ensure a fair comparison.
@@ 5.4 Manipulation Generalization Experiment
Diverse training data is crucial for the generalization of VLA models (Shi et al., 2025). GigaBrain-0 leverages GigaWorld to generate diverse training data, which significantly enhances its robustness and generalization under variations in appearance, object placement, and camera viewpoint. Below, we present our experimental setup and analysis in detail.
@@ 5.5 On-device Experiment
Recent VLA models (Bjorck et al., 2025; Cheang et al., 2025; Zhai et al., 2025; Black et al., 2024; Intelligence et al., 2025; Jiang et al., 2025a), while powerful, are often hindered by their large parameter counts and high computational complexity, making them impractical for deployment on resource-constrained edge devices. However, efficient on-device execution is crucial for real-time robotic applications that require low latency, privacy, and autonomy. To address this challenge, we present GigaBrain-0-Small, an optimized lightweight variant specifically designed for efficient inference on edge platforms such as the NVIDIA Jetson AGX Orin.
Compared to GigaBrain-0, GigaBrain-0-Small adopts the compact vision-language model SmolVLM2 (Marafioti et al., 2025) and reduces the action expert parameters to approximately 100M. Beyond architectural simplification, we implement a series of system-level optimizations: (1) we eliminate redundant CPU–GPU memory transfers and unnecessary data type (dtype) conversions; (2) we enable automatic mixed-precision inference via torch.autocast; (3) we optimize the Rotary Position Embedding (RoPE) (Su et al., 2024) computation by precomputing and caching the sine and cosine lookup tables; and (4) we apply torch.compile to key components, including the denoising step and VLM forward pass, to convert dynamic PyTorch code into optimized static graphs.
@@ 6 Conclusion and Future Work
7 Contributors
The GigaBrain-0 project is the result of collaborative efforts by a dedicated team of researchers and engineers. The contributors, listed below in alphabetical order:
Angen Ye, Boyuan Wang, Chaojun Ni, Guan Huang, Guosheng Zhao, Haoyun Li, Jie Li, Jiagang Zhu, Lv Feng, Peng Li, Qiuping Deng, Runqi Ouyang, Wenkang Qin, Xinze Chen, Xiaofeng Wang, Yang Wang, Yifan Li, Yilong Li, Yiran Ding, Yuan Xu, Yun Ye, Yukun Zhou, Zhehao Dong, Zhenan Wang, Zhichao Liu, Zheng Zhu.

### v3

@@ 4.2 World-Model-Generated Data
Generated Data Quality Inspection. Generated videos inevitably contain hallucinations or artifacts that may degrade downstream training performance. To mitigate this, GigaWorld introduces a comprehensive quality assessment pipeline that evaluates generated videos across multiple dimensions: geometric consistency (robotransfer), multiview consistency (robotransfer), text-description alignment (cosmosreason), and physical plausibility (cosmosreason). Each video is assigned a composite quality score, which determines whether it is suitable for pre-training, fine-tuning, or should be discarded. All model and training details mentioned above will be fully elaborated in the upcoming GigaWorld technical report.
@@ 5.1 Dexterous Manipulation Experiment
Experiment Setup. We evaluate GigaBrain-0 on two dexterous manipulation tasks, laundry folding and paper towel preparation. For laundry folding, we deploy the model on the G1 robot platform and fine-tune it using 300 human-collected demonstration trajectories, the training is conducted with a batch size of 128 for 40K steps. For paper towel preparation, we use 100 demonstrations from PiPER arms, and the training is conducted with a batch size of 128 for 20K steps. We compare GigaBrain-0 against $\pi_{0}$ (pi0), which is implemented using the official open-source code, and fine-tuned with the same training config for fair comparison.
@@ 5.2 Long-horizon Experiment
Experiment Setup. We evaluate GigaBrain-0 on two long-horizon manipulation tasks: table bussing and juice preparation. For table bussing, we deploy the model on the dual-arm PiPER robot platform and fine-tune it using 100 human-collected demonstration trajectories, and the training is conducted with a batch size of 128 for 20K steps. For juice preparation, we utilize the AgiBot G1 robot with 489 demonstrations, and the training is conducted with a batch size of 128 for 35K steps. We compare GigaBrain-0 against $\pi_{0}$ (pi0), which is implemented using the official open-source code, and fine-tuned with the same training config to ensure a fair comparison.
@@ 5.3 Mobile Manipulation Experiment
Experiment Setup. We evaluate GigaBrain-0 on two mobile manipulation tasks: boxes moving and laundry baskets moving. For boxes moving, we deploy the model on the AgiBot G1 robot platform and fine-tune it using 300 human-collected demonstration trajectories, and the training is conducted with a batch size of 128 for 30K steps. For laundry baskets moving, we utilize the dual-arm PiPER robot platform with 378 demonstrations, and the training is conducted with a batch size of 192 for 30K steps. We compare GigaBrain-0 against $\pi_{0}$ (pi0), which is implemented using the official open-source code, and fine-tuned with the same training config to ensure a fair comparison.
@@ 5.4 Manipulation Generalization Experiment
Diverse training data is crucial for the generalization of VLA models (shi2025diversity). GigaBrain-0 leverages GigaWorld to generate diverse training data, which significantly enhances its robustness and generalization under variations in appearance, object placement, and camera viewpoint. Below, we present our experimental setup and analysis in detail.
@@ 5.5 On-device Experiment
Recent VLA models (gr00t; gr3; walloss; pi0; pi05; galaxea), while powerful, are often hindered by their large parameter counts and high computational complexity, making them impractical for deployment on resource-constrained edge devices. However, efficient on-device execution is crucial for real-time robotic applications that require low latency, privacy, and autonomy. To address this challenge, we present GigaBrain-0-Small, an optimized lightweight variant specifically designed for efficient inference on edge platforms such as the NVIDIA Jetson AGX Orin.
Compared to GigaBrain-0, GigaBrain-0-Small adopts the compact vision-language model SmolVLM2 (smolvlm) and reduces the action expert parameters to approximately 100M. Beyond architectural simplification, we implement a series of system-level optimizations: (1) we eliminate redundant CPU–GPU memory transfers and unnecessary data type (dtype) conversions; (2) we enable automatic mixed-precision inference via torch.autocast; (3) we optimize the Rotary Position Embedding (RoPE) (su2024roformer) computation by precomputing and caching the sine and cosine lookup tables; and (4) we apply torch.compile to key components, including the denoising step and VLM forward pass, to convert dynamic PyTorch code into optimized static graphs.